# LaF-GRPO: In-Situ Navigation Instruction Generation for the Visually Impaired via GRPO with LLM-as-Follower Reward

### v1

@@ Abstract
Navigation instruction generation for visually impaired (VI) individuals (NIG-VI) is critical yet relatively underexplored. This study, hence, focuses on producing precise, in-situ, step-by-step navigation instructions that are practically usable by VI users. Concretely, we propose LaF-GRPO (LLM-as-Follower GRPO), where an LLM simulates VI user responses to generate rewards guiding the Vision-Language Model (VLM) post-training. This enhances instruction usability while reducing costly real-world data needs. To facilitate training and testing, we introduce NIG4VI, a 27k-sample open-sourced benchmark. It provides diverse navigation scenarios with accurate spatial coordinates, supporting detailed, open-ended in-situ instruction generation. Experiments on NIG4VI show the effectiveness of LaF-GRPO by quantitative metrics (e.g., Zero-(LaF-GRPO) boosts BLEU +14%; SFT+(LaF-GRPO) METEOR 0.542 vs. GPT-4o’s 0.323) and yields more intuitive, safer instructions. Code and benchmark are available at https://github.com/YiyiyiZhao/NIG4VI.
LaF-GRPO: In-Situ Navigation Instruction Generation for the Visually Impaired via GRPO with LLM-as-Follower Reward
Yi Zhao, Siqi Wang  and Jing Li Department of Computing, The Hong Kong Polytechnic University {yi-yi-yi.zhao, siqi23.wang}@connect.polyu.edu.hk, jing-amelia.li@polyu.edu.hk
@@ 1 Introduction
The Visually Impaired (VI) community, comprising approximately 2.2 billion individuals globally with partial or complete blindness, underscores the significant need for effective assistive technologies. [^1] https://www.who.int/publications/i/item/9789241516570 Enhancing their quality of life through Visually Impaired Assistance (VIA) has motivated extensive research, e.g., VIALM (Zhao et al., 2024) and WalkVLM (Yuan et al., 2025). This paper specifically focuses on Navigation Instruction Generation for VI users (NIG-VI), a VIA sub-area. Navigation Instruction Generation (NIG) was initially conceptualized for general embodied agents, producing high-level trajectory plans. Yet, NIG-VI, being people-centered, operates under fundamentally different constraints. As shown in Figure 1, effective NIG-VI systems must generate in-situ step-level instructions that (1) integrate non-visual sensory cues (auditory/tactile landmarks, surface textures) (2) provide accurate directional and distance guidance to compensate for lack of visual referencing, and (3) adapt to urban obstacles - all within map-coordinate systems while ensuring walking safety.
Figure 1: A NIG4VI sample. (a) The bird’s-eye view map with waypoints and corresponding front-view images. (b) The system takes front-view images, task descriptions, and position data (in the blue box) as input to generate navigation instructions with the benchmark output (in the yellow box) and our model output (in the green box). The grey box indicates a non-VI navigation.
@@
Early attempts such as ASSISTER (Huang et al., 2022) laid the initial groundwork in this field, yet suffered from architectural limitations of BERT-based systems (Devlin et al., 2019). The advent of Vision-Language Models (VLMs) introduces new opportunities through their multimodal understanding and generation capabilities. RL-based post-training methods like GRPO (DeepSeek-AI, 2025) further enhance reasoning abilities, enabling VLMs to align with NIG-VI that demands people-centered guidance. However, most existing approaches rely on large-scale parallel data for fine-tuning, which can be costly and fail to incorporate interactive user feedback essential for people-centered guidance.
@@
To bridge this gap, we propose LLM-as-Follower GRPO (LaF-GRPO)—the first GRPO-based framework for the NIG-VI task, featuring two novel components: (1) an LLM that simulates VI user responses to navigation instructions by interpreting their likely actions, and (2) a VLM post-training procedure for instruction generation, guided by LLM-as-Follower reward. LaF-GRPO mitigates the need for costly VI user trials while ensuring instruction usability with human-in-the-loop navigation simulation. Also, viewing the scarcity of VI navigation benchmarks, we construct NIG4VI - a comprehensive VI navigation instruction benchmark featuring 27k samples for simulation experiments. Fully open-sourced with granular spatial metadata, NIG4VI enables the generation of detailed, open-ended in-situ instructions.
@@
We then experiment with LaF-GRPO on NIG4VI and find: (1) Qwen2.5-VL models trained with Zero-(LaF-GRPO) outperforms the Zero-Shot baseline, achieving superior scores across diverse metrics. (2) Qwen2.5-VL-7B models trained with SFT+(LaF-GRPO) show leading performance, achieving a METEOR score of 0.542, substantially higher than GPT-4o. (3) Beyond quantitative gains, LaF-GRPO can potentially help generate people-centered instructions with enhanced linguistic variety, more intuitive directional cues, richer environmental details, and crucial safety considerations.
In summary, our main contributions are:
@@
$\bullet$ We propose the LaF-GRPO framework, the first method to employ GRPO for NIG-VI with a novel LLM-simulated follower feedback;
@@
$\bullet$ We contribute the NIG4VI benchmark, the first open-source comprehensive dataset featuring precise multi-modal navigation contexts to facilitate robust model evaluation for VI navigation;
@@
$\bullet$ We present extensive empirical studies across VLMs under various paradigms (Zero-shot, Zero-(LaF-GRPO), SFT, and SFT+(LaF-GRPO)), demonstrating the effectiveness of our approach.
@@ 2 Related Work
2.1 VLMs and VIA
@@
VLMs (Liu et al., 2023; Dai et al., 2023; OpenAI, 2024a; Anthropic, 2024; Team, 2024) have drawn attention for combining visual perception with language generation. Refining VLMs with Reinforcement Learning (Ouyang et al., 2022) improves alignment with human preferences and enhances reasoning abilities. Recent success in Group Relative Policy Optimization (GRPO) (DeepSeek-AI, 2025) has led to RL fine-tuned VLMs like VLM-R1 (Shen et al., 2025), AlphaDrive (Jiang et al., 2025), and MedVLM-R1 (Pan et al., 2025), broadening their application range. VIA with VLMs is closely related to visual captioning and Visual Question Answering (VQA). VIALM (Zhao et al., 2024) frames VIA as a VQA task, generating step-by-step guidance from environment images and user requests. While VIALM emphasizes environment-grounded guidance with tactile information, it is not specifically designed for navigation. WalkVLM (Yuan et al., 2025) extends this to dynamic walking assistance and introduces the Walking Awareness Dataset (WAD). Though WalkVLM tackles navigation, its focus remains on video captioning rather than precise orientation and mobility guidance.
2.2 Navigation Instruction Generation (NIG)
There are two main branches for NIG studies: NIG for embodied agents and NIG for the visually impaired. Prior research on NIG for embodied agents has predominantly focused on advanced visual processing techniques while generating trajectory-level instructions. More details can be found in Appendix A. For the NIG-VI branch, ASSISTER (Huang et al., 2022) introduced the UrbanWalk benchmark for in-situ instructional guidance and developed a navigation assistance model. Our work improves upon ASSISTER in two key ways: (1) we introduce a more detailed evaluation benchmark covering orientation, mobility, scene description, and safety warnings; and (2) we leverages advanced VLMs within a GRPO framework with an LLM-as-Follower reward mechanism, leading to more effective navigation instructions.
3 NIG-VI Task and LaF-GRPO Method
@@
Figure 2: Method Overview. Top left: Training sample with input, target output, and generated navigation instruction’s action interpretation. Bottom left: Action interpreter training using LLaMA-3-8B-Instruct to simulate VI users’ navigation responses. Right: Post-training procedures for VLMs processing with LaF-GRPO using multiple reward functions (format, text generation, and LLM-as-Follower reward).
3.1 NIG-VI Task Formulation
@@
We start this section by describing the NIG-VI task of generating in-situ step-by-step natural language instructions to guide VI users along a pre-planned route $P=[p_{1},\dots,p_{K}]$ using a VLM-based assistant system. The route $P$ consists of positional waypoints $p_{i}$ leading to a destination and is generated using the A* algorithm. At each discrete step $i$ of the navigation, the VLM receives two primary inputs: a front-view camera image $x_{\text{image}}^{(i)}$ and a task description which includes the user’s current pose $x_{\text{pose}}^{(i)}=(x_{\text{loc}}^{(i)},x_{\text{rot}}^{(i)})$ represented by their location $x_{\text{loc}}^{(i)}\in\mathbb{R}^{3}$ and rotation $x_{\text{rot}}^{(i)}\in\mathbb{R}^{3}$ within a global map coordinate system, as well as the next target waypoint $p_{i+1}\in P$. Based on these inputs, the VLM $\pi$ generates a sequence of tokens $y=y^{(i)}=(y_{1}^{(i)},y_{2}^{(i)},\dots,y_{t}^{(i)})$ of token length $t$. The generated instruction $y$ might also include details about the current surroundings captured in $x_{\text{image}}^{(i)}$ and any necessary safety alerts:
@@
3.2 The LaF-GRPO Framework
We then discuss our LaF-GRPO framework to tackle the NIG-VI task. It aims to address the challenges of ensuring that navigation instructions are people-centered, practically usable by the VI users, while mitigating the need for costly real-world data collection with VI participants. The overview of LaF-GRPO is illustrated in Figure 2, where the framework comprises two key components: (1) an LLM (without a visual encoder to “see”) that simulates VI users’ responses to navigation instructions by interpreting how these users would act upon hearing the instructions, and (2) a VLM post-training procedure that generates these instructions with (1)’s feedback. LaF-GRPO first employs an action interpreter to produce structured interpretations of potential user actions, which are then used to compute the LLM-as-Follower reward. This reward signal subsequently guides the VLM training process of navigation instruction generation that is more likely to be effectively followed by VI users in real-world navigation scenarios.
Action Interpreter.
The action interpreter models VI user responses to navigation instructions. We fine-tune an LLM $\rho$ with parameters $\varphi$ to predict potential user actions from verbal guidance. Given VLM-generated instruction tokens $y$, it produces a structured action interpretation $\mathcal{A}$. Formally, we define $\mathcal{A}$ as a structured dictionary containing: (1) a ‘move’ action with associated ‘direction’ (indicated using clock positions) and ‘distance’ parameters, and (2) a ‘detailed_hazard_alert’ boolean flag that indicates whether the user perceives warnings about nearby obstacles, as illustrated in Figure 2 Left. To train such an action interpreter, we utilize training samples generated based on the prompt template detailed in Table 8 in Appendix E.
Navigation Instruction Generator.
For VI guidance, we use a pre-trained VLM $\pi$ with parameters $\theta$ for in-situ navigation instruction generation. The training of this generator involves two stages: Supervised Fine-tuning (SFT) and Group Relative Policy Optimization (GRPO). For SFT training details, please refer to Appendix B. For the GRPO, specifically, we propose LaF-GRPO, which is based on the standard GRPO (see below) reward function yet incorporates a novel LLM-as-Follower reward.
@@ 3.2 The LaF-GRPO Framework
Here, the term $\mathcal{L}_{i}$ represents the clipped surrogate objective used in PPO (Schulman et al., 2017):
@@
where $w_{i}=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$ is the importance sampling ratio, $A_{i}$ is the estimated advantage for the output $o_{i}$, based on relative rewards of the outputs inside each group only, calculated as $A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\ldots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\ldots,r_{G}\})}$, and $\epsilon$ is a clipping hyperparameter. The second term, $-\beta\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})$, regularizes the policy by penalizing divergence from a reference policy $\pi_{\text{ref}}$ with coefficient $\beta$. This regularization stabilizes training by keeping the model close to the original effective policy, preventing it from losing previously learned capabilities.
3.3 LaF-GRPO Reward Functions
GRPO leverages verifiable rewards to simplify the reward modeling process. To integrate LLM’s feedback to enable smooth, people-centered language guidance in NIG-VI, LaF-GRPO utilizes three reward functions as follows. The reward calculation algorithm is detailed in Algorithm 1 (Appendix C).
Format Reward.
[…]
Text Generation Reward.
We adopt METEOR to verify text generation rewards to align the generation to the ground-truth style. Here, METEOR is selected based on its evaluation of semantic overlap, incorporating synonymy and stemming to provide a nuanced, human-correlated assessment.
LLM-as-Follower Reward.
To incorporate LLM’s feedback for navigation, we propose the reward $r_{LaF}$ to assess the navigational quality of generated instructions by comparing their interpreted actions (move direction, move distance, and alert flag) against those of a reference. Our intuition is that spatial factors, such as directional accuracy ($a_{dir}$) and movement distance precision ($a_{dist}$), play a direct and critical role in determining navigation success. In contrast, safety alert flags ($a_{alert}$) serve as supplementary support for VI navigation by indicating potential hazards, though they are not primary determinants of success (Giudice and Legge, 2008; Younis et al., 2019). Considering these, the reward is computed as:
@@ 3.3 LaF-GRPO Reward Functions
$\delta(\cdot)$ denotes exact match comparison. To prioritize spatial factors, weighting coefficients are set such that $w_{dir}+w_{dist}>w_{alert}$. This, in the end, yields an $r_{LaF}$ score ranging from 0 to 1.

### v2

@@ Abstract
Navigation instruction generation for visually impaired (VI) individuals (NIG-VI) is critical yet relatively underexplored. This study focuses on generating precise, in-situ, step-by-step navigation instructions that are practically usable for VI users. Specifically, we propose LaF-GRPO (LLM-as-Follower GRPO), where an LLM simulates VI user responses to navigation instructions, thereby providing feedback rewards to guide the post-training of a Vision-Language Model (VLM). This enhances instruction accuracy and usability while reducing costly real-world data collection needs. To address the scarcity of dedicated benchmarks in this field, we introduce NIG4VI, a 27k-sample open-source dataset to facilitate training and evaluation. It comprises diverse navigation scenarios with accurate spatial coordinates, supporting detailed and open-ended in-situ instruction generation. Experiments on NIG4VI demonstrate the effectiveness of LaF-GRPO through quantitative metrics (e.g., Zero-(LaF-GRPO) boosts BLEU 14%; SFT+(LaF-GRPO) METEOR 0.542 vs. GPT-4o 0.323), and qualitative analysis further confirms that our method yields more intuitive and safer instructions.
Codehttps://github.com/YiyiyiZhao/NIG4VI
@@ 1 Introduction
Figure 1: A NIG4VI data sample. (a) Bird’s-eye view map with waypoints and corresponding front-view images. (b) A comparison of instructions generated by ASSISTER, the ground truth, and our model (LaF-GRPO).
@@
The Visually Impaired (VI) community, comprising approximately 2.2 billion (WHO2019vision) individuals globally with partial or complete blindness, underscores the significant need for effective assistive technologies. Extensive research in Visually Impaired Assistance (VIA) has emerged to address this need (zhao2024vialmsurveybenchmarkvisually; yuan2025walkvlmaidvisuallyimpairedpeople; gao2025wearable; cai2024navigating). This paper targets a critical and foundational sub-area: Navigation Instruction Generation for the Visually Impaired (NIG-VI). Unlike Navigation Instruction Generation (NIG) for general embodied agents which produces high-level trajectory plans (DBLP:conf/naacl/DouP22; DBLP:conf/acl/GopinathanMAS24; DBLP:conf/eccv/FanLWY24; DBLP:conf/eccv/KongCWSHYL24), NIG-VI is fundamentally human-centered. As illustrated in Figure 1, it demands the generation of in-situ, step-level instructions that must: (1) integrate non-visual sensory cues like auditory/tactile landmarks and surface textures; (2) provide precise directional and distance guidance to compensate for the lack of visual referencing; and (3) dynamically adapt to urban obstacles, all while ensuring user safety within a map-coordinate system.
@@
Early attempts, such as ASSISTER (DBLP:conf/eccv/HuangSZBBO22), laid the initial groundwork in this field but were ultimately constrained by the architectural limitations of BERT-based systems (devlin-etal-2019-bert). The advent of Vision-Language Models (VLMs) has introduced new opportunities due to their multimodal understanding and generation capabilities. Reinforcement Learning (RL) based post-training methods like GRPO (deepseekai2025deepseekr1incentivizingreasoningcapability) further enhance reasoning abilities, enabling VLMs to align with human preferences, i.e., the human-centered guidance demanded by NIG-VI. However, this alignment process requires collecting large-scale human feedback data for fine-tuning, which can be costly and often fails to incorporate the interactive user feedback essential for achieving human-centered guidance.
@@
To bridge this gap, we propose LLM-as-Follower GRPO (LaF-GRPO), the first GRPO-based framework for the NIG-VI task. It features two novel components: (1) an LLM that simulates how VI users respond to navigation instructions by interpreting their likely actions, and (2) a VLM post-training procedure for instruction generation guided by an LLM-as-Follower reward. LaF-GRPO mitigates the need for costly trials with VI users while ensuring instruction usability through a human-in-the-loop navigation simulation. Furthermore, to address the scarcity of VI navigation benchmarks, we introduce NIG4VI—a comprehensive VI navigation instruction benchmark featuring 27k samples for simulation experiments. Fully open-sourced and annotated with granular spatial metadata, NIG4VI enables the generation of detailed, open-ended, in-situ instructions.
@@
Evaluation of LaF-GRPO on the NIG4VI benchmark reveals the following findings: (1) Qwen2.5-VL models under the Zero-(LaF-GRPO) paradigm significantly outperform the standard zero-shot baseline, demonstrating superior performance across a range of metrics. (2) Qwen2.5-VL-7B models trained with SFT+(LaF-GRPO) show leading performance, achieving a METEOR score of 0.542, substantially surpassing powerful proprietary models like GPT-4o. (3) Beyond quantitative improvements, qualitative analysis further confirms that LaF-GRPO generates more human-centered instructions, distinguished by enhanced linguistic variety, more intuitive directional cues, richer environmental details, and crucial safety considerations. In summary, our main contributions are:
@@
$\bullet$ The LaF-GRPO framework, the first to employ GRPO for NIG-VI with a LLM-simulated follower feedback.
@@
$\bullet$ The NIG4VI benchmark, the first open-source comprehensive dataset with precise multi-modal navigation contexts for robust model evaluation for VI navigation.
@@
$\bullet$ Extensive empirical studies across VLMs under various paradigms (Zero-shot, Zero-(LaF-GRPO), SFT, and SFT+(LaF-GRPO)), validating the method’s effectiveness.
@@ 2 Related Work
VLMs (DBLP:conf/nips/LiuLWL23a; DBLP:conf/nips/Dai0LTZW0FH23; openai2024gpt4technicalreport; TheC3; geminiteam2024gemini15unlockingmultimodal) have gained attention for combining visual perception with language generation. Refining VLMs with Reinforcement Learning (DBLP:conf/nips/Ouyang0JAWMZASR22) improves alignment with human preferences. Recent success in Group Relative Policy Optimization (GRPO) (deepseekai2025deepseekr1incentivizingreasoningcapability) has led to RL fine-tuned VLMs like AlphaDrive (jiang2025alphadriveunleashingpowervlms), VLM-R1 (shen2025vlmr1stablegeneralizabler1style), Praxis-VLM (hu2025praxisvlmvisiongroundeddecisionmaking), and MedVLM-R1 (pan2025medvlmr1incentivizingmedicalreasoning), broadening their applications.
@@
Visually Impaired Assistance (VIA) is a broad and diverse field (gao2025wearable; cai2024navigating; DBLP:conf/chi/GuanX024; DBLP:conf/chi/LiLKC24). VIA with VLMs is closely related to visual captioning and Visual Question Answering (VQA). VIALM (zhao2024vialmsurveybenchmarkvisually) frames VIA as a VQA task, generating guidance from environment images and user requests. While VIALM emphasizes environment-grounded guidance with tactile information, it is not specifically designed for navigation. WalkVLM (yuan2025walkvlmaidvisuallyimpairedpeople) extends this to dynamic walking assistance and introduces the Walking Awareness Dataset (WAD). Though WalkVLM tackles navigation, its focus remains on video captioning rather than precise orientation and mobility guidance.
There are two main branches for NIG studies: NIG for embodied agents and NIG for the visually impaired. Prior research on Navigation Instruction Generation (NIG) for embodied agents has primarily focused on visual processing while generating trajectory-level instructions. BEVInstructor (DBLP:conf/eccv/FanLWY24) employs a Bird’s-Eye View encoder. SAS (DBLP:conf/acl/GopinathanMAS24) uses semantic knowledge with adversarial reward learning. C-Instructor (DBLP:conf/eccv/KongCWSHYL24) focuses on style-controlled instruction generation. Our proposed approach differs in two significant ways: (1) it prioritizes navigation feedback for VLM fine-tuning; and (2) it generates step-level in-situ instructions.
In the NIG-VI field, ASSISTER (DBLP:conf/eccv/HuangSZBBO22) introduced the UrbanWalk benchmark and developed a navigation assistance model. Our work offers two improvements: (1) we introduce a more detailed evaluation benchmark covering orientation, mobility, scene description, and safety warnings, informed by formative studies from the Human–Computer Interaction field that investigate the needs of VI users (merchant2024generatingcontextuallyrelevantnavigationinstructions; zhao2025lessmorereducingcognitive; DBLP:conf/uist/ChangLG24); and (2) we leverage advanced VLMs within a GRPO framework, leading to more effective instructions.
3 Methodology
Grounded in the Speaker-Follower paradigm (DBLP:conf/nips/FriedHCRAMBSKD18), the rationale for LaF-GRPO is to apply Theory-of-Mind (ToM) principles (DBLP:conf/acl/ZhaoND23). Our LLM-as-Follower simulates a user’s cognitive mapping, generating feedback from their anticipated interpretation.
@@ 3 Methodology
Figure 2: Method Overview. Top left: An training sample with input, target output, and action interpretation. Bottom left: Action interpreter training using LLaMA-3-8B-Instruct to simulate VI users’ navigation responses. Right: Post-training for VLMs with LaF-GRPO. The LaF reward differentiates outputs when Format and Text Generation rewards are identical.
NIG-VI Task Formulation
@@
In the NIG-VI task, a VLM-based assistant generates in-situ, step-by-step natural language instructions to guide a visually impaired (VI) user along a pre-planned route $P=[p_{1},\dots,p_{K}]$. The route $P$ consists of positional waypoints $p_{i}$ leading to a destination and is generated using the A* algorithm (DBLP:conf/eccv/HuangSZBBO22). At each discrete step $i$ of the navigation, the VLM receives two primary inputs: a front-view camera image $x_{\text{image}}^{(i)}$ and a task description which includes the user’s current pose $x_{\text{pose}}^{(i)}=(x_{\text{loc}}^{(i)},x_{\text{rot}}^{(i)})$ represented by their location $x_{\text{loc}}^{(i)}\in\mathbb{R}^{3}$ and rotation $x_{\text{rot}}^{(i)}\in\mathbb{R}^{3}$ within a global map coordinate system, as well as the next target waypoint $p_{i+1}\in P$. Based on these inputs, the VLM $\pi$ generates a sequence of tokens $y=y^{(i)}=(y_{1}^{(i)},y_{2}^{(i)},\dots,y_{t}^{(i)})$ of token length $t$. The generated instruction $y$ might also include details about the current surroundings captured in $x_{\text{image}}^{(i)}$ and any necessary safety alerts:
@@
The LaF-GRPO Framework
We propose LaF-GRPO to address the challenges of generating navigation instructions that are human-centered and practically usable for VI users, while mitigating the need for costly real-world data collection with VI participants. The overview of LaF-GRPO is illustrated in Figure 2, where the framework comprises two key components: (1) an action interpreter LLM (without a visual encoder to ‘see’) that simulates VI users’ responses to navigation instructions by interpreting how these users would act upon hearing the instructions, and (2) a VLM post-training procedure that generates these instructions with (1)’s feedback.
Action Interpreter. To simulate a VI user’s response to navigation instructions, we use Supervised Fine-tuning (SFT) to train an LLM $\rho_{\varphi}$ to predict potential actions. Prior to its deployment, the interpreter is validated to have a parse accuracy above 98% on a held-out set, ensuring its reliability for generating reward signals. Given VLM-generated instruction tokens $y$, it produces a structured action interpretation $\mathcal{A}$. Formally, we define $\mathcal{A}$ as a structured dictionary containing: (1) a ‘move’ action with associated ‘direction’ (indicated using clock positions (yuan2025walkvlmaidvisuallyimpairedpeople)) and ‘distance’ parameters, and (2) a ‘detailed_hazard_alert’ boolean flag that indicates whether the user perceives warnings about nearby obstacles, as illustrated in Figure 2 Left.
Navigation Instruction Generator. For VI guidance, we use a pre-trained VLM $\pi$ with parameters $\theta$ for in-situ navigation instruction generation. The training of this generator involves two stages: Supervised Fine-tuning (SFT) and Group Relative Policy Optimization (GRPO). Our proposed LaF-GRPO framework enhances standard GRPO (see below) by incorporating a novel LLM-as-Follower reward.
@@
Here, the term $\mathcal{L}_{i}$ represents the clipped surrogate objective used in PPO (schulman2017proximalpolicyoptimizationalgorithms):
@@
where $w_{i}=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$ is the importance sampling ratio, $A_{i}$ is the estimated advantage for the output $o_{i}$, based on relative rewards of the outputs inside each group only, calculated as $A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\ldots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\ldots,r_{G}\})}$, and $\epsilon$ is a clipping hyperparameter. The second term, $\beta\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})$, regularizes the policy by penalizing divergence from a reference policy $\pi_{\text{ref}}$ with coefficient $\beta$. This regularization stabilizes training by keeping the model close to the original effective policy, preventing it from losing previously learned capabilities.
LaF-GRPO Reward Functions
LaF-GRPO utilizes three reward functions as follows.
Format Reward. To encourage controllable generation, we adopt this binary reward ($r_{\text{format}}\in\{0,1\}$) that evaluates structural compliance with the expected response format. Here, the reward would equal 1 if the output follows the required format pattern $think>\n$<answer>.*?</answer>’ in sequence, and 0 otherwise.
Text Generation Reward. The text generation reward ($r_{meteor}$) is calculated as the METEOR score between the output and the ground-truth reference. METEOR is selected based on its evaluation of semantic overlap, incorporating synonymy and stemming to provide a nuanced assessment.
LLM-as-Follower Reward. The LLM-as-Follower reward ($r_{LaF}$) assesses the navigational quality of generated instructions by comparing their interpreted actions (move direction, distance, and alert flag) with those of a ground-truth reference. The rationale behind this design is that spatial factors, such as directional accuracy ($a_{dir}$) and movement distance precision ($a_{dist}$), play a direct and critical role in determining navigation success. In addition, safety alert flags ($a_{alert}$) serve as supplementary support for VI navigation by indicating potential hazards, though they are not primary determinants of success (giudice2008blind; younis2019smart). Accordingly, we formulate the reward as:
@@
$\delta(\cdot)$ denotes an exact match. $r_{LaF}$ is in the range [0, 1].